# Video Depth Anything: Consistent Depth Estimation for Super-Long Videos

### v2

@@ 1 Introduction
Specifically, we first design a lightweight spatial-temporal head (STH) to replace the DPT head [28] and enable temporal information interactions. STH includes four temporal attention layers, applied along the temporal dimension for each spatial position. Introducing temporal attention only in the head prevents the learned representation from being corrupted by the limited video data. Then, we propose a temporal gradient matching loss to constrain depth prediction gradients along the temporal dimension to match those calculated from the ground truth. This loss function is jointly optimized with the scale-shift-invariant loss and spatial gradient matching loss [3, 41]. Despite its simplicity, it can effectively boost the model’s temporal consistency. Third, to maintain the original capabilities of the model, we train it jointly on 550K video frames with depth annotations using supervised learning, and on 0.62 million unlabeled images using self-training, similar to Depth Anything V2 [42]. To handle super-long videos at inference time, we developed a novel segment-wise processing strategy. Each new segment is concatenated with eight overlapping frames and two key frames from the previous video clips, forming a total of 32 frames. Then, the overlapping frames will be progressively interpolated between the two consecutive windows to ensure smoothness.
@@ 3.2 Temporal Gradient Matching loss
[…]
@@ 3.3 Inference strategy for super-long sequence
Figure 3: Inference strategy for long videos. $N$ is the video clip lenght consumed by our model. Each inference video clip is built by $N-T_{o}-T_{k}$ future frames, $T_{o}$ overlapping/adjacent frames, and $T_{k}$ key frames. The key frames are selected by taking every $\Delta_{k}$-th frame going backward. Then, the new depth predictions will be scale-shift-aligned to the previous frames based on the $T_{k}$ overlapping frames. We use $N=32,T_{o}=8,T_{k}=2,\Delta_{k}=12$.
@@
Key-frame referencing. As illustrated in  Fig.˜3, a subsequent video clip for inference is composed of three parts: $N-T_{o}-T_{k}$ future frames, $T_{o}$ overlapping frames from the previous clip and $T_{k}$ key frames. The key frames are subsampled from the previous frames with an interval of size $\Delta_{k}$. Therefore, the video clip to be consumed share the same length as during training. This approach incorporates content from earlier windows into the current window with minimal computation burden. According to our experiment results, such simple strategy can significantly reduce accumulated scale drift, especially for long video.
@@ 4 Experiments
DAv2-L [42] 0.137 0.815 0.150 0.768 0.127 0.864 0.094 0.928 0.390 0.541 1.140 3.6
NVDS [38] 0.233 0.614 0.207 0.628 0.199 0.674 0.217 0.598 0.408 0.464 2.176 6.8
NVDS [38] + DAv2-L [42] 0.227 0.617 0.194 0.658 0.191 0.700 0.184 0.679 0.449 0.503 2.536 5.8
ChoronDepth [32] 0.243 0.576 0.199 0.665 0.199 0.665 0.173 0.771 0.192 0.673 1.022 5.2
DepthCrafter [13] 0.164 0.753 0.169 0.730 0.153 0.803 0.141 0.822 0.299 0.695 0.639 3.4
@@
VDA-S (Ours) 0.086 0.942 0.110 0.876 0.083 0.950 0.077 0.959 0.339 0.584 0.703 2.6
VDA-L (Ours) 0.083 0.944 0.089 0.926 0.071 0.959 0.062 0.971 0.295 0.644 0.570 1.6
Table 1: Zero-shot video depth estimation results. We compare with representative single-image [42] and video depth estimation models [38, 32, 13, 40]. “VDA-S” denotes our model with ViT-Small backbone. “VDA-L” denotes our model with ViT-Large backbone. The best and the second best results are highlighted.
@@ 4.2 Zero-shot Depth Estimation
[…]
Image depth results. As displayed in Table 2, our video depth model achieves competitive depth metrics compared to DAv2-L [42] in most datasets. This shows that our model maintains the geometric accuracy of the foundation model while also ensuring video stabilization.
@@
[…]
@@
Table 3: Inference latency comparisons for video depth estimation. We measure average runtime for each frame on a single A100 GPU with a resolution of $\times 518$ 518.
@@ 3 Limitations and future work
Our model is trained on publicly available video depth datasets without specially curating data, which may limit its capabilities due to the data quantity. We believe that with more data, the model’s performance can be further improved, and the backbone network can be unlocked for fine-tuning. Additionally, although our model is significantly more computationally efficient than the baselines, it still faces challenges in handling streaming videos, which we leave as future work.
@@ 4 More Details of Pipeline
The weights are initialized from Depth Anything V2 [42]. Training comprises two stages. In the first stage, synthetic and wild binocular data are used. In the second stage, synthetic videos and unlabeled single images are employed. Besides the loss defined in Equation 4 used for synthetic videos, unlabeled single images are supervised using the same method as described in [42]. During training, we uniformly sample video clips of 32 frames from each dataset, resize the shorter edge of images to 518 pixels, and perform random center cropping, resulting in training clips with a resolution of 518×518×32518$\times$ 518$\times$ 32518 × 518 × 32. We use the AdamW [20] optimizer with a cosine scheduler, setting the base learning rate to $1e^{-4}$. The batch size is set to 16 for video frames, each with a length of 32 frames, and 128 for image datasets. The loss weights for the single frame loss, TGM loss, and distillation loss are set to 1.0, 10.0, and 0.5, respectively.
@@ 5 More Details of Evaluation
[…]
@@ 6 Applications
By aligning single frame with metric depth, which can be obtained from a metric depth model or a sparse point cloud acquired through SLAM, our model can generate a depth point cloud for the entire environment using camera information. The generated point cloud can then be transformed into a mesh and utilized for 3D reconstruction, AR, and VR applications. We present a point cloud generation case in Fig. 12. Here, we sample 10 frames spanning approximately 5 seconds from the KITTI dataset [11]. After obtaining the inferred inverse depth, we compute the global scale and shift by aligning the first frame with the corresponding metric inverse depth. We then apply the affine transformation to the entire set of inverse depth frames and convert them to depth. The final point cloud is generated by merging the point clouds from each frame. As shown in Fig. 12, our model generates a clean and regular point cloud compared to DepthCrafter [13] and Depth Any Video [40].

### v3

@@ 1 Introduction
Specifically, we first design a lightweight spatial-temporal head (STH) to replace the DPT head [28] and enable temporal information interactions. STH includes four temporal attention layers, applied along the temporal dimension for each spatial position. Introducing temporal attention only in the head prevents the learned representation from being corrupted by the limited video data. Then, we propose a temporal gradient matching loss to constrain depth prediction gradients along the temporal dimension to match those calculated from the ground truth. This loss function is jointly optimized with the scale-shift-invariant loss and spatial gradient matching loss [3, 42]. Despite its simplicity, it can effectively boost the model’s temporal consistency. Third, to maintain the original capabilities of the model, we train it jointly on 730K video frames with depth annotations using supervised learning, and on 0.62 million unlabeled images using self-training, similar to Depth Anything V2 [43]. To handle super-long videos at inference time, we developed a novel segment-wise processing strategy. Each new segment is concatenated with eight overlapping frames and two key frames from the previous video clips, forming a total of 32 frames. Then, the overlapping frames will be progressively interpolated between the two consecutive windows to ensure smoothness.
@@ 3.2 Temporal Gradient Matching loss
Here, $d_{i},g_{i}$ are scaled and shifted versions of the predictions and ground truth. $\hat{d_{i}},\hat{g_{i}}$ denotes the warped depth from the subsequent frame using optical flow. $\mid\cdot\mid$ is used to represent the absolute values.
@@ 3.3 Inference strategy for super-long sequence
Figure 3: Inference strategy for long videos. NN is the video clip lenght consumed by our model. Each inference video clip is built by $N-T_{o}-T_{k}$ future frames, $T_{o}$ overlapping/adjacent frames, and $T_{k}$ key frames. The key frames are selected by taking every $\Delta_{k}$-th frame going backward. Then, the new depth predictions will be scale-shift-aligned to the previous frames based on the $T_{k}$ overlapping frames. We use $N=32,T_{o}=8,T_{k}=2,\Delta_{k}=12$.
@@
Key-frame referencing. As illustrated in  Fig.˜3, a subsequent video clip for inference is composed of three parts: $N-T_{o}-T_{k}$ future frames, $T_{o}$ overlapping frames from the previous clip and $T_{k}$ key frames. The key frames are subsampled from the previous frames with an interval of size $\Delta_{k}$. Therefore, the video clip to be consumed share the same length as during training. This approach incorporates content from earlier windows into the current window with minimal computation burden. Furthermore, we carefully select the values of $T_{k}$ and $\Delta_{k}$ to ensure that the first frame of a video is always positioned at the beginning of each clip, thereby enhancing depth consistency for extended videos. According to our experiment results, such simple strategy can significantly reduce accumulated scale drift, especially for long video.
@@ 4 Experiments
DAv2-L [43] 0.137 0.815 0.150 0.768 0.127 0.864 0.094 0.928 0.390 0.541 1.140 5.5
NVDS [39] 0.233 0.614 0.207 0.628 0.199 0.674 0.217 0.598 0.408 0.464 2.176 8.5
NVDS [39] + DAv2-L [43] 0.227 0.617 0.194 0.658 0.191 0.700 0.184 0.679 0.449 0.503 2.536 7.7
ChoronDepth [32] 0.243 0.576 0.199 0.665 0.199 0.665 0.173 0.771 0.192 0.673 1.022 6.6
DepthCrafter [13] 0.164 0.753 0.169 0.730 0.153 0.803 0.141 0.822 0.299 0.695 0.639 5.0
@@
VDA-S (Ours-Syn) 0.089 0.937 0.124 0.839 0.081 0.955 0.087 0.953 0.326 0.596 0.702 4.0
VDA-L (Ours-Syn) 0.083 0.946 0.087 0.933 0.070 0.961 0.064 0.967 0.300 0.633 0.570 1.7
VDA-S (Ours) 0.086 0.942 0.110 0.876 0.083 0.950 0.077 0.959 0.339 0.584 0.703 3.9
VDA-L (Ours) 0.083 0.944 0.089 0.926 0.071 0.959 0.062 0.971 0.295 0.644 0.570 1.7
Table 1: Zero-shot video depth estimation results. We compare with representative single-image [43] and video depth estimation models [39, 32, 13, 41]. “VDA-S” and “VDA-L” denote our model with ViT-Small and ViT-Large backbones, respectively. Models denoted with “-Syn” are trained exclusively on public datasets, as detailed in supplementary. The best and the second best results are highlighted.
@@ 4.2 Zero-shot Depth Estimation
Video depth results. As demonstrated in Table 1, our VDA model achieves state-of-the-art performance across all long video datasets, excelling in both geometric and temporal metrics. This underscores the effectiveness of our robust foundation model and the innovative design of our video model. Notably, on the KITTI [11], Scannet [7], and Bonn [24] datasets, our model surpasses other leading methods by a significant margin of approximately $10\%$ in the geometric accuracy metric $\delta_{1}$, although it is trained on much fewer video data compared to DepthCrafter [13] (over 10 million frames) and DepthAnyVideo [41] (6 million frames). For the short video synthetic dataset Sintel [5], where sequences contain around 50 frames each, DepthCrafter [13] exhibits better accuracy than our model. This discrepancy may be attributed to the absence of movie data, which features frames with focal lengths similar to those in Sintel [5], in our model’s training set. It is also worth highlighting that our compact model, VDA-S, which has significantly lower latency compared to other models (as shown in Table 3), demonstrates superior geometric accuracy over representative diffusion-based methods for long videos.
Image depth results. The datasets in Table 2 consist of single images, identical to those utilized in DAv2 [43]. Although our model is designed for videos, it demonstrates the capability to handle static images. As shown in Table 2, our video depth model achieves competitive depth metrics compared to DAv2-L [43] across most datasets. This demonstrates that our model not only retains the geometric accuracy of the foundational model but also delivers high accuracy and consistency in video depth estimation.
@@
Inference time. We measure the inference latency of various models on an A100 GPU. As shown in Table 3, our large model achieves the lowest inference time compared to both diffusion-based methods (DepthAnyVideo [41] and DepthCrafter [13]) and the transformer-based method (NVDS [39]). This performance is attributed to our feed-forward transformer structure and lightweight temporal modules. Notably, the latency of our large model, VDA-L, is only approximately $10\%$ greater than that of DAv2-L [43, 13], which uses the same encoder structure, thus demonstrating the efficiency of our spatiotemporal head. Furthermore, the inference latency of our small model is less than 10ms, indicating its potential for real-time applications.
@@
Table 3: Inference latency comparisons for video depth estimation. We measure average runtime for each frame on a single A100 GPU with a resolution of $518\times 518$.
@@ 3 Limitations and future work
Our model is trained primarily on publicly available video depth datasets, which may limit its capabilities due to the data quantity. We believe that with more data, the model’s performance can be further improved, and the backbone network can be unlocked for fine-tuning. Additionally, although our model is significantly more computationally efficient than the baselines, it still faces challenges in handling streaming videos, which we leave as future work.
@@ 4 More Details of Pipeline
The weights are initialized from Depth Anything V2 [43]. Training comprises two stages. In the first stage, synthetic and wild binocular data are used. In the second stage, synthetic videos and unlabeled single images are employed. Models labeled ’-Syn’ in the main paper are exceptionally trained using synthetic videos and unlabeled images in a single stage. Besides the loss defined in Equation 4 of the main paper used for synthetic videos, unlabeled single images are supervised using the same method as described in [43]. During training, we uniformly sample video clips of 32 frames from each dataset, resize the shorter edge of images to 518 pixels, and perform random center cropping, resulting in training clips with a resolution of $518\times 518\times 32$. We use the AdamW [20] optimizer with a cosine scheduler, setting the base learning rate to $1e^{-4}$. The batch size is set to 16 for video frames, each with a length of 32 frames, and 128 for image datasets. The loss weights for the single frame loss, TGM loss, and distillation loss are set to 1.0, 10.0, and 0.5, respectively.
@@ 5 More Details of Evaluation
Here, $f$ represents the projection function that maps the depth $\hat{x}^{k}_{d}$ from the $k$-th frame to the $(k+1)$-th frame using the transformation matrix $p^{k}$. $p^{k+1}_{-}$ is the inverse matrix for inverse projection. $N$ denotes the number of frames.
@@ 6 Applications
By aligning single frame with metric depth, which can be obtained from a metric depth model or a sparse point cloud acquired through SLAM, our model can generate a depth point cloud for the entire environment using camera information. The generated point cloud can then be transformed into a mesh and utilized for 3D reconstruction, AR, and VR applications. We present a point cloud generation case in Fig. 12. Here, we sample 10 frames spanning approximately 5 seconds from the KITTI dataset [11]. After obtaining the inferred inverse depth, we compute the global scale and shift by aligning the first frame with the corresponding metric inverse depth. We then apply the affine transformation to the entire set of inverse depth frames and convert them to depth. The final point cloud is generated by merging the point clouds from each frame. As shown in Fig. 12, our model generates a clean and regular point cloud compared to DepthCrafter [13] and Depth Any Video [41]. Point cloud generation for wild videos is illustrated in Fig. 13. Compared to DepthCrafter [13] and DepthAnyVideo [41], our model produces more regular point clouds.
@@
Figure 13: Point cloud generation for wild videos. We compare our method with DepthCrafter [13] and DepthAnyVideo [41] using three videos from DAVIS dataset [26]. Camera intrinsics, along with aligned scale and shift parameters, are derived from processing the first frame of each video through MoGe [36]. Point cloud distortions are highlighted with red boxes.